A block slides down a 100m inclined plane. Suppose that the incline is frictionless, what is the speed of the block as it reaches the bottom of the incline? 


1
Expert's answer
2021-02-07T19:08:22-0500

We can find the final velocity of the block at the bottom of the incline from the law of conservation of energy:


"PE=KE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot100\ m}=44.27\ \dfrac{m}{s}."